# SCMM: Calibrating Cross-modal Representations for Text-Based Person Search

### v5

@@ IV-D2 Impact of Sew Calibration Loss for Reducing Cross-modal Gap
IV-D3 Impact of Masking Caption Modeling
The masking caption modeling operation in fine-grained cross-modal interaction is critical in our framework. It comprises tokens masking, attention module, and masked tokens prediction. As shown in Table V, we explore the effectiveness of MCM components on the CUHK-PEDES. First, only masking on the input text tokens without a reconstruction task behaves like a random erase text augmentation. This augmentation can already bring +1.36% marginal improvement and reach 63.65% on Rank1. It shows that the details provided by word tokens are helpful for fine-grained recognition.
TABLE VI: Performance analysis for cross-domain validation.
@@ IV-D3 Impact of Masking Caption Modeling
IV-D4 Analysis of Generalisability Validation
To better understand the contributions of MCM in our framework, we conduct a domain generalization analysis on the three benchmarks to demonstrate our generalization performance. In Table VI, CUHK $\Rightarrow$ ICFG and CUHK $\Rightarrow$ RSTP indicate using the model trained on the CUHK-PEDES dataset to infer their test sets and vice versa. Compared to the baseline, we can observe a performance improvement in the three metrics for ICFG-PEDES and RSTPReid. The fine-grained features mined by MCM are more resistant to overfitting. The improvement in domain generalization shows the capability of our framework in learning generic cross-modal information, which is essential to solve the fine-grained modal heterogeneity.
IV-D5 Effect of Margin Parameters
In subsection  III-A, we introduce the adaptive margins to constrain cross-modal learning using Sew Calibration loss with constraints. We find that compared to the manually fixed parameters (SEW-f), our adaptive margins (SEW-a) produce better results as shown in Fig. 6, where we investigate the effect of manual margin in different settings on the CUHK-PEDES [10] on Rank1 accuracy. The best performance of 65.41% is achieved when $\mathcal{M}=0.5$, while the worst model performance of 62.29% is observed when $\mathcal{M}=0$. We note that the performance of the model first improves as the margin increases. This is due to our bi-directional margin compressing the same pedestrian target compactly. However, we find that the model performance does not continue to improve as the margin increases, and we observe an inflection point around margin 0.5. When the margin is too large, the tight constraints make the model too hard to train.
@@ IV-D5 Effect of Margin Parameters
IV-D6 Effect of Mask Ratio
Inspired by Masked Language Modeling [37], we propose the MCM loss which leverages a masked captions prediction task to establish detailed and generic relationships between textual and visual parts. Our method requires setting the mask ratio in the MCM. To investigate the effect of the mask ratio on model performance and generalization learning ability, we conducted ablation experiments, and the results are shown in Table VII. When the mask ratio is set to 0, which is equivalent to the baseline, the performance is the worst. Compared to the baseline, our method with a 0.1 mask ratio achieves the best results, with scores of 65.16%, 81.95%, and 87.93% on Rank1, Rank5, and Rank10, respectively.
TABLE VII: Ablation results of mask ratio on the CUHK-PEDES.
@@ IV-D6 Effect of Mask Ratio
The experiments are conducted on the CUHK-PEDES dataset using the Baseline+MCM method. As shown in these results, we find that the model achieves the best performance when the mask is set to 0.1, and then the model performance gradually decreases as the mask ratio increases. This is because the information provided in the annotated caption text is very accurate and free of redundant information. Therefore, as the mask ratio expands, more and more key information may be obscured, causing the model to fail to learn enough semantic information to complete the construction of the image-text relationship.
@@
IV-E Qualitative Analysis
IV-E1 Visualization of Attention Map
@@
IV-E2 Visualization of Image-text Feature Distribution
@@
SCMM is capable of learning better image-text representation distributions with fine-grained image-text interaction and adaptive constraints to reduce the cross-modal gap. To demonstrate this, we randomly selected some pedestrians and extracted their image-text global features, then mapped them to two dimensions for visualization in Fig. 7(a) and (b). For instance, we can take IDs 5532 and 11800 as examples. We can observe that the image-text features of the same ID distribution are compressed more compactly, and the boundaries between them are more apparent than the baseline results. This demonstrates the effectiveness of SCMM in mitigating the cross-modal gap.
V Conclusion
In this work, we introduce SCMM, a novel method for TBPS that addresses the critical need for precise and efficient person retrieval in IoT environments. SCMM enhances cross-modal representation learning to reduce the inter-modal gap and discern fine-grained image-text correspondences. This is achieved through a dual-pronged strategy employing Sew calibration and MCM losses, which together refine feature alignment and enrich cross-modal correspondences. The method’s streamlined architecture ensures high-speed inference, essential for real-time applications in IoT settings. For instance, SCMM can be integrated with edge computing units in surveillance cameras to preprocess video streams and extract textual descriptions of individuals, facilitating efficient person search across the network. This integration is crucial for applications in smart cities and industrial environments, where timely and accurate person identification is vital for maintaining security and operational efficiency. Extensive experiments on three popular TBPS benchmarks demonstrate the effectiveness and superiority of SCMM. Future work will explore enhancing SCMM’s generalizability to multi-modal inputs, investigating its potential in multi-view TBPS scenarios, and addressing challenges in domain adaptation and few-shot learning, thereby advancing TBPS towards more practical and versatile applications in smart surveillance and multimedia ecosystems.

### v7

@@ 4.4.3 Impact of Masked Caption Modeling
4.4.4 Analysis of Generalisability Validation
Table 6: Performance analysis for cross-domain validation.
@@ 4.4.4 Analysis of Generalisability Validation
To better understand the contributions of MCM in our framework, we conduct a domain generalization analysis on the three benchmarks to demonstrate our generalization performance. In Table 6, CUHK $\Rightarrow$ ICFG and CUHK $\Rightarrow$ RSTP indicate using the model trained on the CUHK-PEDES dataset to infer their test sets and vice versa. Compared to the baseline, we can observe a performance improvement in the three metrics for ICFG-PEDES and RSTPReid. The fine-grained features mined by MCM are more resistant to overfitting. The improvement in domain generalization shows the capability of our framework in learning generic cross-modal information, which is essential to solve the fine-grained modal heterogeneity.
4.4.5 Effect of Margin Parameters
In Sec.  3.1, we introduce the adaptive margins to constrain cross-modal learning using Sew Calibration loss with constraints. We find that compared to the manually fixed parameters (SEW-f), our adaptive margins (SEW-a) produce better results as shown in Fig. 5, where we investigate the effect of manual margin in different settings on the CUHK-PEDES [8] on Rank1 accuracy. The best performance of 65.41% is achieved when $\mathcal{M}=0.5$, while the worst model performance of 62.29% is observed when $\mathcal{M}=0$. We note that the performance of the model first improves as the margin increases, which is due to our bi-directional margin compressing the same pedestrian target compactly. However, we find that the model performance does not continue to improve as the margin increases, and we observe an inflection point around margin 0.5. When the margin is too large, the tight constraints make the model too hard to train.
@@ 4.4.5 Effect of Margin Parameters
4.4.6 Effect of Mask Ratio
Inspired by Masked Language Modeling [36], we propose the MCM loss which leverages a masked captions prediction task to establish detailed and generic relationships between textual and visual parts. Our method requires setting the mask ratio in the MCM. To investigate the effect of the mask ratio on model performance and generalization learning ability, we conducted ablation experiments, and the results are shown in Table 7. When the mask ratio is set to 0, which is equivalent to the baseline, the performance is the worst. Compared to the baseline, our method with a 0.1 mask ratio achieves the best results, with scores of 65.16%, 81.95%, and 87.93% on Rank1, Rank5, and Rank10, respectively.
Table 7: Ablation results of mask ratio on the CUHK-PEDES.
@@ 4.4.6 Effect of Mask Ratio
The experiments are conducted on the CUHK-PEDES dataset using the Baseline+MCM method. As shown in these results, we find that the model achieves the best performance when the mask is set to 0.1, and then the model performance gradually decreases as the mask ratio increases. The main reason is that the information provided in the annotated caption text is very accurate and free of redundant information. Therefore, as the mask ratio expands, more and more key information may be obscured, causing the model to fail to learn enough semantic information to complete the construction of the image-text relationship.
@@
4.5 Qualitative Analysis
4.5.1 Visualization of Attention Map
@@
Figure 8: Top 5 retrieval results on CUHK-PEDES. Matched target person images are marked with red rectangles. and subcaptions show the text queries used for retrieval.
4.5.2 Visualization of Image-text Feature Distribution
@@
SCMM is capable of learning better image-text representation distributions with fine-grained image-text interaction and adaptive constraints to reduce the cross-modal gap. To demonstrate this, we randomly selected some pedestrians and extracted their image-text global features, then mapped them to two dimensions for visualization in Fig. 7(a) and (b). For instance, we can take IDs 5532 and 11800 as examples. We can observe that the image-text features of the same ID distribution are compressed more compactly, and the boundaries between them are more apparent than the baseline results, which demonstrates the effectiveness of SCMM in mitigating the cross-modal gap.
4.5.3 Visualization of Top 5 Search Qualitative Results
To evaluate the performance of our method, we randomly selected several sets of search results from our method and visualized them in Fig. 8. For each target image, we show the top 5 most similar search results, with the similarity scores after re-ranking processing displayed above each image. Correct search results are indicated by red rectangles, and the number in the last column shows how many pedestrian targets with the same ID have been recalled. Taking ID 5219 for an example, our method correctly retrieved all three target images, with similarity scores of 1.2563, 1.2100, and 0.9775, respectively. For the last two unmatched images, our method gave similarity scores of 0.5174 and 0.5011, respectively. Although these images have some similarities with the positive samples, they do not match the text description in other details, and our method learned to differentiate them. The last row in Fig. 8 shows a hard case, where the man in the query image has extremely similar attributes to several other pedestrians. Despite, our method is able to correctly rank 2/3 of the right person images as the top 2 search results.
5 Conclusion
In this work, we introduce SCMM, a novel AI framework for text-based person search that addresses fundamental challenges in cross-modal learning. Our method introduces two innovative components: sew calibration loss with adaptive margins and masked caption modeling loss that establishes fine-grained correspondences through transformer-based architectures. These complementary strategies bridge the inter-modal gap while capturing detailed relationships essential for person identification. Extensive experimental validation demonstrates SCMM’s superior performance, achieving state-of-the-art results of 73.81%, 64.25%, and 57.35% Rank1 accuracy on CUHK-PEDES, ICFG-PEDES, and RSTPReID respectively. SCMM’s streamlined dual-encoder architecture prioritizes computational efficiency, making it suitable for real-time applications. Future work will explore incorporating additional modalities and developing more sophisticated adaptive mechanisms for multimodal learning in person re-identification tasks.
CRediT authorship contribution statement
Jing Liu: Conceptualization, Methodology, Formal analysis, Writing – original draft, Writing – review & editing, Visualization. Donglai Wei: Methodology, Software, Validation, Formal analysis, Investigation, Writing – review, Project administration & editing. Yang Liu: Conceptualization, Methodology, Investigation, Writing – review & editing, Visualization. Sipeng Zhang: Writing – review & editing, Funding acquisition. Tong Yang: Writing – review & editing, Funding acquisition. Wei Zhou: Conceptualization, Writing – review & editing. Weiping Ding: Methodology, Writing – review & editing. Victor C. M. Leung: Supervision, Writing – review & editing, Project administration.
Data availability
Data will be made available on request.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used generative AI and AI-assisted technologies for writing enhancement and language polishing to improve the clarity and readability of the manuscript. After using these tools, the authors have reviewed and edited the content as needed and take full responsibility for the content of the publication.
Acknowledgements
This work was supported in part by the Mitacs Elevate under Grant No. IT44479; the Guangdong Pearl River Talents Recruitment Program under Grant No. 2019ZT08X603; and the Guangdong Pearl Rivers Talent Plan under Grant No. 2019JC01X235. We sincerely thank all the editors and anonymous reviewers for their careful work and thoughtful suggestions that have helped improve this paper substantially.